# Six-Degree-of-Freedom Aircraft Landing Trajectory Planning with Runway Alignment

### v1

@@ Abstract
$S$, $S_{t}$  = platform area for the tail and the wing
@@ 1 Introduction
Optimization-based trajectory generation is a promising direction because it can provide a systematic framework to enforce state and control constraints, including the equations of motions (aircraft dynamics), i.e., it can generate dynamically feasible trajectories [2]. In addition to imposing constraints, another motivation for using an optimization-based approach is its capability in optimizing key mission objectives such as minimizing time and fuel. Due to these advantages, optimization-based approaches have been extensively utilized for various applications: powered descent guidance for reusable rockets [3, 4, 5, 6, 7], hypersonic reentry [8, 9, 10], path generation and tracking for ground vehicles [11, 12], control of humanoid robots [13, 14], and spacecraft rendezvous [15, 16].
@@ 4.1 Runway alignment constraint
The position results are given in Figure 4, with a detailed zoom-in of the runway provided in Figure 5. The results of the total cost (61) and the iteration count are summarized in Table 4. All trajectories are generated with the identical parameters and tolerances as detailed in Table 3 with the weights $w_{vc}=10^{2}$ and $w_{tr}=1$ and the extrapolation parameter $\gamma=1$, with the exception of the trajectory computed using the STC method starting point (B). The velocity and Euler angle profiles are shown in Figure 6, and the angular velocity and control input profiles are illustrated in Figure 7 and Figure 8, respectively.
It is clear from Figure 4 and Figure 5 that the trajectories computed using the proposed method are aligned with the runway around the touch down point, as intended, and hence they are more operational. Meanwhile the trajectories computed without the runway alignment constraint tend to either be too steep or too shallow during the final landing phase, neither of which are reasonable in practice. The cost function values for the cases without the alignment constraint are smaller on average than those for the cases with the alignment constraint. This is expected since a trajectory with fewer constraints can be expected to have a lower cost. The trajectories computed using the standard STC method are aligned with the runway as well, but they tend to have larger iteration counts. This is because, as discussed before, the STC constraint is implemented as a nonconvex constraint (27), thereby adding the nonconvexity of the problem. Furthermore, we can see in Figure 5 that the trajectories computed using the standard STC method exhibit large inter-sample constraint violations with respect to the alignment constraint. One remedy for this is to combine our method for the continuous-time constraint satisfaction given in Section 3.6 with the STC constraint, but this combination will add the nonconvexity of the problem even further.
@@
Figure 5: Close-up views of the runway from Figure 4. The pentahedron represented by dash-dotted green lines is the feasible region with respect to the runway alignment constraint. The first column from the left shows the result of the proposed method, the second column shows that of the standard STC method, and the third column shows the case without the runway alignment constraint. For each case, the second and the third plots from the top show the position results projected onto North-Up and West-Up planes, respectively. Circular marks on trajectories illustrate node points.
@@ 4.4 Monte Carlo study
Next, we re-examine the failed cases by repeating the previously described procedure, but this time, starting with weight parameters $w_{vc}=10^{3}$ and $w_{tr}=10$. For $\gamma=1$, all 9 out of the 9 failed cases are resolved; for $\gamma=1.2$, 9 out of the 10 failed cases are resolved; for $\gamma=1.4$, 9 out of the 11 failed cases are resolved; and for $\gamma=1.6$, 15 out of the 46 failed cases are corrected. As a result, the proposed algorithm achieves success rates of 100% for $\gamma=1.0$, 99.67% for $\gamma=1.2$, 99.33% for $\gamma=1.4$, and 89.67% for $\gamma=1.6$. In conclusion, the results of the Monte Carlo study show that the proposed method converges reliably for a wide range of initial conditions, and that the xPTR algorithm can accelerate the PTR algorithm, especially with extrapolation parameter values between 1.0 and 1.4.
@@ Acknowledgments
We thank the members of the Autonomous Controls Laboratory (ACL) at the University of Washington, especially Purnanand Elango for the discussion and valuable feedbacks. We also thank the members of Boeing, especially Dragos Margineantu, for their valuable guidance and helpful comments.

### v2

@@ Abstract
$S$, $S_{t}$  = platform area for the wing and the tail
@@ 1 Introduction
Optimization-based trajectory generation is a promising direction because it can provide a systematic framework to enforce state and control constraints, including the equations of motion (aircraft dynamics), i.e., it can generate dynamically feasible trajectories [2]. In addition to imposing constraints, another motivation for using an optimization-based approach is its capability in optimizing key mission objectives such as minimizing time and fuel. Due to these advantages, optimization-based approaches have been extensively utilized for various applications: powered descent guidance for reusable rockets [3, 4, 5, 6, 7], hypersonic reentry [8, 9, 10], path generation and tracking for ground vehicles [11, 12], control of humanoid robots [13, 14], and spacecraft rendezvous [15, 16].
@@ 3.6 Extrapolation update
3.7 Convergence analysis
Without the extrapolation step ($\gamma=1$), the solution method aligns with the framework of the prox-linear method [65, 47], enabling the exploitation of its convergence guarantees. The prox-linear method is designed to solve
$\Theta({\bf y})=J({\bf y})+H(G({\bf y})),$
where ${\bf y}$ represents the decision variable. Here, $H$ is a convex, $\alpha_{h}$-Lipschitz continuous function, and $G$ is potentially nonconvex continuously differentiable with a $\beta_{g}$-Lipschitz continuous gradient. To optimize $\Theta({\bf y})$, the prox-linear method iteratively minimizes a convex approximation of $\Theta({\bf y})$, given by
$\Theta^{\rho}({\bf y};{\bf y}_{i})=J({\bf y})+H(G({\bf y}_{i})+\nabla G({\bf y}_{i})({\bf y}-{\bf y}_{i}))+\frac{1}{2\rho}\lVert{\bf y}-{\bf y}_{i}\rVert_{2}^{2},$
where the subscript $i$ denotes the current iteration index. In our specific setup, ${\bf y}$ is a concatenated vector of all decision variables, expressed as
${\bf y}=({\bf x}_{1},{\bf u}_{1},s_{1},\ldots,{\bf x}_{N-1},{\bf u}_{N-1},s_{N-1},{\bf x}_{N},{\bf u}_{N}),$
and the constant $\rho$ is given by $1/(2w_{tr})$. The terms $J({\bf y})$, $H(G({\bf y}))$, and $G({\bf y})$ are
$J({\bf y})$ $=w_{t}J_{\text{time}}^{d}+w_{\mathrm{thr}}J_{\mathrm{thr}}^{d}+w_{r}J_{r}^{d}+w_{\Omega}J_{\Omega}^{d}+I_{\mathcal{Y}}({\bf y}),$
$H(G({\bf y}))$ $=w_{vc}J_{vc},$
$G({\bf y})$ $=({\bf x}_{1}-{\bf x}_{0}-\lim_{\tau\rightarrow 1}\int_{0}^{\tau}s_{1}{\bf f}({\bf x}(s),{\bf u}(s))\text{d}s,\cdots,{\bf x}_{N}-{\bf x}_{N-1}-\lim_{\tau\rightarrow 1}\int_{0}^{\tau}s_{N-1}{\bf f}({\bf x}(s),{\bf u}(s))\text{d}s,$
$\quad\ \ \int_{t_{0}}^{t_{1}}\max\{{\bf h}({\bf x}(t)),0\}^{2}\text{d}t,\cdots,\quad\ \ \int_{t_{N-1}}^{t_{N}}\max\{{\bf h}({\bf x}(t)),0\}^{2}\text{d}t),$
where $I_{\mathcal{Y}}$ is an indicator function, and $\mathcal{Y}$ represents a feasible set to convex constraints (40)-(43). The prox-linear method has the following convergence result.
Theorem 1
(Lemma 5.1 of [65]) At iteration $i$ of the prox-linear method, the following holds
$\Theta({\bf y}_{i})\geq\Theta({\bf y}_{i+1})+\frac{\rho}{2}(2-\alpha_{h}\beta_{g}\rho)\lVert\mathcal{G}_{\rho}({\bf y}_{i})\rVert^{2},$
where $\mathcal{G}_{\rho}(\cdot)$ is the prox-gradient mapping defined as
$\mathcal{G}_{\rho}({\bf y})=\frac{1}{\rho}\left(z-\operatorname*{argmin}_{{\bf y}^{\prime}}\Theta^{\rho}({\bf y}^{\prime};{\bf y})\right).$
Hence, $\Theta^{\rho}({\bf y})$ is monotonically decreasing for $\rho\leq\frac{1}{\alpha_{h}\beta_{g}}$.
[…]
@@ 4.1 Runway alignment constraint
The position results are given in Figure 4, with a detailed zoom-in of the runway provided in Figure 5. In Figure 5, the pentahedron represented by dash-dotted green lines is the feasible region with respect to the runway alignment constraint. The first column from the left shows the result of the proposed method, the second column shows that of the standard STC method, and the third column shows the case without the runway alignment constraint. For each case, the second and the third plots from the top show the position results projected onto North-Up and West-Up planes, respectively. Circular marks on trajectories illustrate node points.
The results of the total cost (61) and the iteration count are summarized in Table 4. All trajectories are generated with the identical parameters and tolerances as detailed in Table 3 with the weights $w_{vc}=10^{2}$ and $w_{tr}=1$ and the extrapolation parameter $\gamma=1$, with the exception of the trajectory computed using the STC method starting point (B). The velocity and Euler angle profiles are shown in Figure 6, and the angular velocity and control input profiles are illustrated in Figure 7 and Figure 8, respectively.
It is clear from Figure 4 and Figure 5 that the trajectories computed using the proposed method are aligned with the runway around the touch down point, as intended, and hence they are more operational. Meanwhile the trajectories computed without the runway alignment constraint tend to either be too steep or too shallow during the final landing phase, neither of which are reasonable in practice. The cost function values for the cases without the alignment constraint are smaller on average than those for the cases with the alignment constraint. This is expected since a trajectory with fewer constraints can be expected to have a lower cost. The trajectories computed using the standard STC method are aligned with the runway as well, but they tend to have larger iteration counts. As pointed out previously, this is due to having the STC constraint implemented as a nonconvex constraint (27), thereby making the feasible set more challenging to optimize over. Furthermore, we can see in Figure 5 that the trajectories computed using the standard STC method exhibit large inter-sample constraint violations with respect to the alignment constraint. One remedy for this is to combine the method for the continuous-time constraint satisfaction given in Section 3.6 with the STC constraint, but this combination will further exacerbate the nonconvexity of the problem.
@@
Figure 5: Close-up views of the runway from Figure 4. Further details are provided in the main text.
@@ 4.4 Monte Carlo study
Next, we re-examine the failed cases by repeating the previously described procedure, but this time, starting with weight parameters $w_{vc}=10^{3}$ and $w_{tr}=20$. For $\gamma=1$, all 9 out of the 9 failed cases are resolved; for $\gamma=1.2$, 10 out of the 10 failed cases are resolved; for $\gamma=1.4$, 10 out of the 11 failed cases are resolved; and for $\gamma=1.6$, 36 out of the 46 failed cases are corrected. As a result, the proposed algorithm achieves success rates of 100% for $\gamma=1.0$, 100% for $\gamma=1.2$, 99.67% for $\gamma=1.4$, and 96.67% for $\gamma=1.6$. In conclusion, the results of the Monte Carlo study show that the proposed method converges reliably for a wide range of initial conditions, and that the xPTR algorithm can accelerate the PTR algorithm, especially with extrapolation parameter values between 1.0 and 1.4.
@@ Acknowledgments
This work was supported by Boeing under Grant 2021-PD-PA-471. We thank the members of the Autonomous Controls Laboratory (ACL) at the University of Washington, especially Purnanand Elango for the discussion and valuable feedback. We also thank Dragos Margineantu for valuable guidance and helpful suggestions.